# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### v2

@@ Abstract
Recent approaches for arbitrary-scale single image super-resolution (ASSR) have used local neural fields to represent continuous signals that can be sampled at arbitrary rates. However, the point-wise query of the neural field does not naturally match the point spread function (PSF) of a given pixel, which may cause aliasing in the super-resolved image. We present a novel way to design neural fields such that points can be queried with an adaptive Gaussian PSF, so as to guarantee correct anti-aliasing at any desired output resolution. We achieve this with a novel activation function derived from Fourier theory. Querying points with a Gaussian PSF, compliant with sampling theory, does not incur any additional computational cost in our framework, unlike filtering in the image domain. With its theoretically guaranteed anti-aliasing, our method sets a new state of the art for ASSR, while being more parameter-efficient than previous methods. Notably, even a minimal version of our model still outperforms previous methods in most cases, while adding 2-4 orders of magnitude fewer parameters. Code and pretrained models are available at https://github.com/prs-eth/thera.
@@ 1 Introduction
For years, learning-based image super-resolution (SR) methods have achieved ever better results. But unlike interpolation methods, which allow us to resample a given image at any resolution, those methods require retraining for each desired scaling factor. Recently, several schemes for arbitrary-scale SR (ASSR) have been proposed that allow end-users to choose any scaling factor, rendering the method much more flexible [17]. Notably, with LIIF, Chen et al. [11] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors. Their method has inspired several follow-ups which build upon the idea of using local neural fields [22, 8, 10]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [46]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal strength, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable. Among others, they have been widely adopted for 3D reconstruction with radiance [32], occupancy [31, 37], and distance fields [35, 39].
Figure 1: Given a low-resolution input, a hypernetwork predicts the parameters of a customized neural field $\Phi_{\bm{\theta}}(x,y,t)$ composed of sinusoidal components. We obtain Gaussian-filtered (anti-aliased) versions of the image that correspond to arbitrary pixel sizes by simply modulating the components according to their frequencies (insets illustrate components decaying at different rates). This representation can then be queried at arbitrary coordinates $(x,y,t)$ to recover pixel grids at any desired scaling factor.
Neural fields offer a compelling representation of continuous signals. However, for many tasks, it is not desirable to query such fields at a single point, but rather to have an integral observation model such as a point spread function (PSF) [3]. ASSR calls for a neighborhood-dependent observation model, as the signal should be observed at various sampling frequencies and is prone to aliasing. Incorporating a physically plausible observation model is not trivial [3, 4, 25, 49, 16, 5], but has the potential to avoid aliasing. For this reason, Chen et al. [11] already took a first step towards learning multiscale representations, via cell encoding. For radiance fields, Mip-NeRF [3] and subsequent work [4, 16, 5] have explored ways to mitigate aliasing effects by focusing on the positional encoding of the input coordinates, as a means to control different frequencies in the field. Lindell et al. [25] and Yang et al. [49] have succeeded in performing subband decomposition of neural fields, though bands are architecturally fixed to a small number of discrete intervals and cannot be modified after training.
In our work, we combine recent advances in implicit neural representations with ideas from classical signal theory to develop a neural field that guarantees anti-aliasing by construction. Interestingly, this filtering incurs no additional overhead: The cost of querying field values is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths. The key insight is that sinusoidal activation functions [40] offer the possibility to selectively decay each component according to its spatial frequency, following Fourier theory. Inspired by the connection between Fourier decompositions and the heat equation, we propose Thera, a neural field method for ASSR underpinned by a novel thermal activation function, see Fig. 1. That formulation makes it possible to analytically compute Gaussian-blurred versions of the neural field $\Phi(\mathbf{x})$ for any desired (isotropic) blur radius. I.e., one can query $\Phi$ with a Gaussian PSF that is matched to the output resolution to prevent aliasing. [^1] While our framework is in principle not bound to the heat equation or Gaussian PSFs, it is an intuitive interpretation for readers familiar with signal theory. Thera provides theoretical guarantees w.r.t. anti-aliasing and multi-scale representation, in exchange for certain constraints on the neural architecture that can easily be overcome with a hypernetwork [14]. Empirically, it outperforms competing methods in most of our experiments and is far more parameter-efficient. We demonstrate that a lightweight variant of our model still outperforms most previous methods while adding 2-4 orders of magnitude fewer parameters. To the best of our knowledge, Thera is the first neural field method to allow bandwidth control at test time. Our main contributions are:
@@
    We introduce thermal activations, which enable neural fields with a principled Gaussian observation model, and therefore anti-aliasing with minimal overhead.
@@
    We use thermal activations to build Thera, a novel method for ASSR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
@@ 2 Related Work
2.0.1 Arbitrary-scale SR.
In ASSR the scaling factor can be chosen at inference time to be any positive number, which allows for maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [17], which infers the parameters of a convolutional upsampling layer using a hypernetwork [14] conditioned on the desired scaling factor. An influential successor work is LIIF [11], in which the high-resolution image is implicitly described by local neural fields that are conditioned with features extracted from the low-resolution input image. The continuous nature of the neural field allows for sampling target pixels at arbitrary locations, and thus also arbitrary resolution. Following work has since built upon the LIIF framework. For instance, UltraSR [48] improves modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for neural radiance fields [32, 3, 4, 5]. LTE [22] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [43] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. Most recently, methods have integrated attention mechanisms into their implicit representation, enabling the neural field to adaptively attend to image features depending on the query coordinates and scale factor, improving reconstruction quality [8, 10]. These methods provide the required scaling factor explicitly as an input to the field, thus attempting to learn the observation model from data. In this paper, we explore a way to directly integrate a (Gaussian) observation model into the neural field representation.
@@
Implicit neural representations. A neural field is a neural network trained to map coordinates onto values of some physical quantity. Recently, neural fields have been used for parameterizing various types of visual data, including images [20, 40, 41, 11, 22, 12, 45], 3D scenes (e.g., represented as signed distance fields [35, 40, 39, 44, 45], occupancy fields [31, 37], LiDAR fields [19], view-dependent radiance fields [32, 3, 4, 5, 45]), or digital humans [50, 53, 9, 47, 13]. We refer the interested reader to Xie et al. [46] for a comprehensive survey of neural fields in vision and beyond. Frequently, it is desirable to impose some prior over the space of learnable representations. A common approach for such conditioning is encoder-based inference [46], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [11, 22, 43, 8, 10]. The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [14]. The latter is a trainable neural network $\Psi\colon\bm{z}\mapsto\bm{\theta}$, directly mapping latent codes to parameters of the field $\bm{\theta}$. An early example of this approach, which is gaining popularity [46], was proposed by Sitzmann et al. [40].
@@
Anti-aliasing in neural fields. Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [3] proposed integrating the positional encoding with Gaussian weights, which reduced aliasing in NeRF. Improvements were later proposed for unbounded scenes [4] and to improve efficiency [16]. Barron et al. [5] tackle anti-aliasing for the increasingly popular Instant-NGP [33] approach. Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [25] or polynomial neural fields [49], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks. Such representations would not be a good fit for ASSR since they do not allow for continuous anti-aliasing, nor bandwidth control at test time. In ASSR, methods usually provide the desired scale as input and expect to learn it from data [11, 8].
@@ 3 Method
To expound how anti-aliasing can be computed analytically, we must first define a sampling model and recap a well-established result from Fourier analysis. For simplicity, we stick to 2-dimensional signals $\mathbf{x}\in\mathbb{R}^{2}$, but all theory is valid for any number of dimensions. We refer the reader to [34] for a comprehensive discussion of sampling and aliasing.
To minimize aliasing when sampling a continuous signal $f(\mathbf{x})$ it is necessary to suppress high-frequency parts of the signal that cannot be represented at the desired sampling rate. One must apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$. We will show that one can naturally integrate this behavior into a neural field’s architecture. The key to this is that, if a signal is decomposed into sinusoidal components, filtering can be done simply by re-scaling each component by a frequency-dependent factor. This equivalence is depicted in Fig. 2. Decomposing a given signal into sinusoidal components is normally a costly operation, but it turns out that we can avoid it by designing the neural field such that the decomposition itself is the native representation of the signal.
3.1 Thermal Activation Functions
The intuition behind our method goes back to the starting point of Fourier analysis: the heat equation. Other implicit representation methods, such as SIREN [40], optimize the parameters $\bm{\theta}$ of a field $\Phi(\mathbf{x})$ by minimizing a data fidelity term $\mathbb{E}_{(\mathbf{x}_{i},\mathbf{y}_{i}\sim\mathcal{D})}\left[\mathcal{L}\left(\Phi(\mathbf{x}_{i}),\mathbf{y}_{i}\right)\right]$ over a dataset $\mathcal{D}$. Our key insight is that we can interpret this field as the temperature distribution on an (isotropic) surface at a given time $t_{0}$. It is well known how to solve the heat equation forwards in time – its solution is obtained by blurring the initial state at $t_{0}$ with a Gaussian kernel whose radius depends on the time difference $t-t_{0}$. Our field $\Phi(\mathbf{x},t)$ now has an additional temporal dimension, but given a single slice $\Phi(\mathbf{x},t_{0})$ there is a unique solution $\Phi(\mathbf{x},t)$ for every $t\geq t_{0}$ such that the field obeys the heat equation
$\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi=\kappa\cdot\left(\frac{\partial^{2}\Phi}{\partial x_{1}^{2}}+\frac{\partial^{2}\Phi}{\partial x_{2}^{2}}\right)\;,$ (1)
with $\kappa>0$ the thermal diffusivity of the medium. In contrast, the solution for $0\leq t<t_{0}$ is ill-posed, mirroring the fact that image downsampling is well-posed, but super-resolution is not.
Figure 2: If we have access to the sinusoidal decomposition of a signal, filtering can be done analytically by scaling each component appropriately. We design our neural field so that we have access to such a decomposition at no extra cost.
It turns out that we can design the field’s architecture in such a way that it is guaranteed to respect Eq. 1. We do so by formulating $\Phi$ as a two-layer perceptron
@@ 3.1 Thermal Activation Functions
where $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$ are MLP parameters, $\nu$ refers to the wave numbers (frequencies) implied by $\mathbf{W}_{1}$ (described in more detail later), and the thermal activation function $\xi(\cdot)$ models the decay of different sinusoidal components over time. We set $t=0$ to be the initial thermal state of a given system represented as a sum of sinusoidal components:
$\xi(\mathbf{z},\nu,\kappa,0)=\sin(\mathbf{z}).$ (3)
For $t>0$, each component is decayed according to its wave number, the thermal diffusivity constant $\kappa$, and $t$:
@@
To obtain $\left|\nu(\mathbf{W}_{1})\right|$ we can simply compute the row-wise Euclidean norm of $\mathbf{W}_{1}$, corresponding to the magnitudes of the implied wave numbers (direction is irrelevant, as all components with the same frequency decay equally, regardless of their orientation). Equation 4 results in higher-frequency components decaying faster than lower-frequency ones, giving rise to a Gaussian low-pass filter whose strength increases with time $t$.
There is an ideal bijection between the sampling rate $f_{s}$ and $t$. At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$). A low-pass filtered version of the signal is observed for $t>0$. To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$. The relationship between the cut-off frequency of the filter and $t$ is controlled by $\kappa$, which we can freely set to any positive number. For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$. Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
$\kappa=\frac{\sqrt{\log(4)}}{2f_{s}\pi^{2}}.$ (5)
To subsample the signal $\mathcal{D}$ by a factor $S$, $\Phi$ should be sampled at
@@
The derivations of these values can be found in Appendix 0.A.
In our formulation, $\mathbf{W}_{1}$ can be seen as a frequency bank whose components (“basis functions”) make up the signal $\Phi(\mathbf{x},0)$, and the vector $\mathbf{b}_{1}$ represents the phase shifts of those components. The matrix $\mathbf{W}_{2}$, with one row per output channel, contains the initial ($t=0$) magnitudes of these components, and $\mathbf{b}_{2}$ is the global bias of $\Phi$ per channel.
Image domain
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA $\times$1/32 AA
Neural field
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$ $\Phi(\mathbf{x},32^{2})$
Figure 3: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with thermal activations (bottom row), which is computed without convolutions or over-sampling. The neural field was supervised only at $\Phi(\mathbf{x},1)$.
Figure 3 shows an example where we fit the field at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 6. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
By definition, our field $\Phi(\mathbf{x},t)$ follows the heat equation and is therefore able to perform anti-aliasing at arbitrary sampling frequencies. Note that modulating deeper fields, as is done in [30], would no longer guarantee anti-aliasing. An alternative (equivalent) interpretation is to view $\Phi(\mathbf{x},t)$ as an observation of the true, continuous signal $\Phi(\mathbf{x},0)$ with a Gaussian point spread function whose width (standard deviation) is determined by $t$.
3.2 Thermal Activations for Super-Resolution

### v3

# Thera: Aliasing-Free Arbitrary-Scale Super-Resolution with Neural Heat Fields

@@ Abstract
Recent approaches to arbitrary-scale single image super-resolution (ASR) use neural fields to represent continuous signals that can be sampled at arbitrary resolutions. However, point-wise queries of neural fields do not naturally match the point spread function (PSF) of pixels, which may cause aliasing in the super-resolved image. Existing methods attempt to mitigate this by approximating an integral version of the field at each scaling factor, compromising both fidelity and generalization. In this work, we introduce neural heat fields, a novel neural field formulation that inherently models a physically exact PSF. Our formulation enables analytically correct anti-aliasing at any desired output resolution, and – unlike supersampling – at no additional cost. Building on this foundation, we propose Thera, an end-to-end ASR method that substantially outperforms existing approaches, while being more parameter-efficient and offering strong theoretical guarantees. The project page is at https://therasr.github.io.
Figure 1: We present Thera, the first method for arbitrary-scale super-resolution with a built-in physical observation model. Given an input image, a hypernetwork predicts the parameters of a specially designed neural heat field, inherently decomposing the image into sinusoidal components. The field’s architecture automatically attenuates frequencies as a function of the scaling factor so as to match the output resolution at which the signal is re-sampled.
@@ 1 Introduction
Over the years, learning-based image super-resolution (SR) methods have achieved increasingly better results. However, unlike interpolation techniques that can resample images at any resolution, these methods typically require retraining for each scaling factor. Recently, arbitrary-scale SR (ASR) approaches have emerged, which allow users to specify any desired scaling factor without retraining, significantly increasing flexibility [18]. Notably, with LIIF, Chen et al. [10] pioneered the use of neural fields for single-image SR, exploiting their continuous representation to enable SR at arbitrary scaling factors. LIIF has since inspired several follow-ups which build upon the idea of using per-pixel neural fields [24, 7, 9, 56]. This is not surprising: Neural fields are in many ways a natural match for variable-resolution computer vision and graphics [48]. By implicitly parameterizing a target signal as a neural network that maps coordinates to signal value, they offer a compact representation, defined over a continuous input domain, and are analytically differentiable.
While neural fields naturally model continuous functions, they do not easily allow for observations of such functions other than point-wise evaluations. For many tasks, however, integral observation models such as point spread functions (PSFs) are desirable. This is particularly true for neural fields-based ASR methods, which by nature do not commit to a fixed upscaling factor a priori but regress continuous representations with unbounded spectra that can be observed at various sampling rates. If the Nyquist frequency corresponding to the desired sampling rate is lower than the highest frequency represented by the field, the sampling operation is prone to aliasing. This explains the initially counterintuitive relevance of anti-aliasing for super-resolution: When using neural fields, signals are first upsampled to infinite (continuous) resolution and then resampled at the desired resolution, and this latter operation must be done carefully. Incorporating a physically plausible observation model is not trivial [2, 3, 27, 51, 17, 4], but has the potential to avoid aliasing. For this reason, Chen et al. [10] and successor works [24, 7, 9, 56] have already taken a first step towards learning multi-scale representations, via cell encoding. Fundamentally, these “learning-based anti-aliasing” approaches try to learn an integrated version of the field for any scaling factor, wasting field capacity for approximating a relation that can be described exactly through Fourier theory.
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$31.531.531.531.531.631.631.631.631.731.731.731.731.831.831.831.831.931.931.931.93232323232.132.132.132.1Thera AirThera PlusThera Pro(Ours)MetaSRLIIFCUFCiaoSRCLITMSITSRNOModel Size (M parameters)PSNR (dB)
Figure 2: Comparison of recent ASR methods, averaged over ×$\times\{2,3,4\}$ scales. We generally achieve higher performance at lower parameter counts. Our best model, Thera Pro, achieves highest overall performance by a large margin.
In this work, we combine recent advances in implicit neural representations with ideas from classical signal theory to introduce neural heat fields, a novel type of neural field that guarantees anti-aliasing by construction. The key insight is that sinusoidal activation functions [41] enable selective attenuation of individual components depending on their spatial frequency, following Fourier theory. This allows for the exact computation of Gaussian-blurred versions of the field for any desired (isotropic) blur radius. When rasterizing an image, the field can therefore be queried with a Gaussian PSF that matches the target resolution, effectively preventing aliasing. Notably, filtering with neural heat fields incurs no computational overhead: The querying cost is the same for any width of the anti-aliasing filter kernel, including infinite and zero widths.
Building on this, we then propose Thera, an end-to-end ASR method that combines a hypernetwork [16] with a grid of local neural heat fields, offering theoretical guarantees with respect to multi-scale representation (see Fig. 1). Empirically, Thera outperforms all competing ASR methods, often by a substantial margin, and is more parameter-efficient (see Fig. 2). To the best of our knowledge, Thera is also the first neural field method to allow bandwidth control at test time.
In summary, our main contributions are:
@@
    We introduce neural heat fields, which represent a signal with a built-in, principled Gaussian observation model, and therefore allow anti-aliasing with minimal overhead.
@@
    We use neural heat fields to build Thera, a novel method for ASR that offers theoretically guaranteed multi-scale capabilities, delivers state-of-the-art performance and is more parameter efficient than prior art.
@@ 2 Related Work
2.1 Neural Fields
@@
A neural field, also called an implicit neural representation, is a neural network trained to map coordinates onto values of some physical quantity. Recently, neural fields have been used for parameterizing various types of visual data, including images [22, 41, 42, 10, 24, 12, 47], 3D scenes (e.g., represented as signed distance fields [36, 41, 40, 46, 47], occupancy fields [33, 37], LiDAR fields [20], view-dependent radiance fields [34, 2, 3, 4, 47]), or digital humans [52, 55, 8, 49, 15]. Frequently, it is desirable to impose some prior over the space of learnable implicit representations. A common approach for such conditioning is encoder-based inference [48], where a parametric encoder maps input observations to a set of latent codes $\bm{z}$, which are often local [10, 24, 44, 7, 9]. The encoded latent variables $\bm{z}$ are then used to condition the neural field, for instance by concatenating $\bm{z}$ to the coordinate inputs or through a more expressive hypernetwork [16], mapping latent codes $\bm{z}$ to neural field parameters $\bm{\theta}$. An early example of this approach, which is gaining popularity [48], was proposed in [41].
2.2 Arbitrary-Scale Super-Resolution
ASR is the sub-field of single-image SR in which the desired SR scaling factor can be chosen at inference time to be (theoretically) any positive number, allowing maximum flexibility, such as that of interpolation methods. The first work along this line is MetaSR [18], which infers the parameters of a convolutional upsampling layer using a hypernetwork [16] conditioned on the desired scaling factor. An influential successor work is LIIF [10], in which the high-resolution image is implicitly described by local neural fields. These fields are conditioned via concatenation, with features extracted from the low-resolution input image. The continuous nature of the neural fields allows for sampling target pixels at arbitrary locations and thus also arbitrary resolution.
Most subsequent work has since been built upon the LIIF framework. For example, UltraSR [50] improves the modeling of high-frequency textures with periodic positional encodings of the coordinate space, as is common practice for e.g., neural radiance fields [34, 2, 3, 4]. LTE [24] makes learning higher frequencies more explicit by effectively implementing a learnable coordinate transformation into 2D Fourier space, prior to a forward pass through an MLP. Vasconcelos et al. [44] use neural fields in CUF to parameterize continuous upsampling filters, which enables arbitrary-scale upsampling. More recently, methods like CiaoSR [7], CLIT [9], and most recently MSIT [56] have integrated (multi-scale) attention mechanisms, improving reconstruction quality. In a parallel line of research, Wei et al. [45] propose SRNO, an attention-based neural operator that learns a continuous mapping between low- and high-resolution function spaces.
2.3 Anti-Aliasing in Neural Fields
@@
Early in the recent development of implicit neural representations, concerns regarding aliasing were raised. Barron et al. [2] proposed integrating a positional encoding with Gaussian weights, which reduced aliasing in NeRF [34]. Improvements were later proposed for unbounded scenes [3] and to improve efficiency [17]. In [4], Barron et al. tackle anti-aliasing within the Instant-NGP [35] approach. Recent work has succeeded in limiting the bandwidth using multiplicative filter networks [27], polynomial neural fields [51] or cascaded training [39], although these works are restricted to discrete, pre-defined band limits (and thus resolutions) and have not tackled super-resolution tasks. These methods are not a good fit for ASR because they do not allow for continuous anti-aliasing, nor bandwidth control at test time. To perform scale-dependent filtering, most fields-based ASR methods instead explicitly provide the scale as input to the field, attempting to learn an appropriate observation model from data. While this approach may work reasonably well in in-distribution settings, it seeks to learn a model from data that can be described exactly with a differential equation, ultimately sacrificing fidelity and generalization.
In contrast, in this paper we explore a way to directly integrate a physics-informed observation model into the neural field representation.
@@ 3 Method
Figure 3: Overview of Thera. A hypernetwork estimates parameters $\{\bm{b}_{1},\bm{W}_{2}\}^{(i,j)}$ of pixel-wise, local neural heat fields. The phase shifts $\bm{b}_{1}$ operate on globally learned components, before thermal activations scale each component depending on their frequency and the desired scaling factor. The components are then linearly combined using coefficients $\bm{W}_{2}$, resulting in an appropriately-blurred, continuous local neural field. This field is then rasterized at the appropriate sampling rate (resolution) to yield a part of the final output image (red square).
In this section we introduce Thera, a novel neural fields-based ASR method that guarantees analytical anti-aliasing at any desired output resolution at no additional cost. First, we present neural heat fields, a special type of neural field that inherently achieves anti-aliasing by implicitly attenuating high-frequency components as a function of a time coordinate. Next, we propose a mechanism for learning a prior over a grid of neural heat fields, enabling them to represent a multi-scale output image conditioned on a lower-resolution input image. Finally, we show that our formulation allows us to impose a regularizer on the underlying, continuous signal itself – something that, to the best of our knowledge is not possible in previous methods.
3.1 Neural Heat Fields for Analytical Anti-Aliasing
Let $\mathbf{x}\in\mathbb{R}^{2}$ denote the spatial coordinates of a continuous image function $f(\mathbf{x})$. Aliasing occurs when this continuous signal is sampled at a rate that does not adequately capture its highest frequency components, resulting in overlapping spectral replicas in the Fourier domain. One must therefore apply a low-pass filter $g(\mathbf{x})$ whose cut-off frequency is aligned with the Nyquist frequency of the sampling rate, then sample the band-limited signal $f\circledast g(\mathbf{x})$. The key of our method is that, if a signal is decomposed into sinusoidal components, such filtering can be done simply by re-scaling each component by a factor that depends on their frequency and a time coordinate $t$. This behavior is naturally accomplished by parameterizing the field $\Phi$ as a two-layer perceptron,
@@ 3.1 Neural Heat Fields for Analytical Anti-Aliasing
with parameters $\bm{\theta}:=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$. Intuitively, $\mathbf{W}_{1}$ serves as a frequency bank, with its components acting as the basis functions that compose the signal $\Phi(\mathbf{x},0)$, and phase shifts encoded by $\mathbf{b}_{1}$. The matrix $\mathbf{W}_{2}$, with one row per output channel, contains initial magnitudes of these components, and $\mathbf{b}_{2}$ is the global bias of $\Phi$ per channel. Finally, we introduce the thermal activation function $\xi(\cdot)$, which models the aforementioned decay of sinusoidal components (implied by $\mathbf{W}_{1}$) over time:
@@
Here, $\left|\nu\right|=\left|\nu(\mathbf{W}_{1})\right|$ denotes the row-wise Euclidean norm of $\mathbf{W}_{1}$, representing the magnitudes of the implied wave numbers (frequencies). Interestingly, Eq. 1 constitutes the solution of the isotropic heat equation $\frac{\partial\Phi}{\partial t}=\kappa\cdot\nabla_{\mathbf{x}}^{2}\Phi$, as derived in Appendix A. We therefore refer to this MLP as a neural heat field.
There is an ideal bijection between the desired sampling rate $f_{s}$ and $t$. At $t=0$ no filtering takes place, implying a continuous signal ($f_{s}\to\infty$). A low-pass filtered version of the signal is observed for $t>0$. To obtain a desired level of anti-aliasing, we only need to compute the corresponding value of $t$. The relationship between the cut-off frequency of the filter and $t$ is controlled by the diffusivity constant $\kappa$, which we can freely set to any positive number. For simplicity, and without loss of generality, we set $\kappa$ such that the native resolution of the (observed, discrete) signal $\mathcal{D}$ is $t=1$. Assuming equal sampling rate $f_{s}$ along both coordinate axes, the optimal value of $\kappa$ then evaluates to
$\kappa=\frac{\sqrt{\ln(4)}}{2f_{s}\pi^{2}}.$ (3)
To subsample the signal $\mathcal{D}$ by a factor $S$, the field $\Phi$ should be sampled at
@@
For a derivation of these values and a demonstration of the filtering mechanism of neural heat fields, see Appendix A.
3.2 Learning a Super-Resolution Prior